C2H5OH + 3O2 -> 2CO3 + 3H20

a) What mass of oxygen is required to react with 1200 grams of ethanol?
b) if 655 grams of water is produced, what mass of ethanol is burned?
1
Expert's answer
2020-01-24T04:59:29-0500

C2H5OH + 3O2"\to" 2CO2 + 3H2O

a) What mass of oxygen is required to react with 1200 g of ethanol

n(C2H5OH) =1200(g)/46(g/mol)=26.09 mol

n(O2)=26.09/1=X/3 n of O2=78.3 mol O2

m(O2)= n*M =78.3*32=2504 g


b) If 655 g of water is produced, what mass of ethanol is burned?

n(H2O) =655g/18(g/mol)=36.39 mol

n of H2O/n of H2O = n of C2H5OH/n of C2H5OH 36.39/3=X/1 n of C2H5OH=12.13 mol m(C2H5OH)=12.13 X (2(12)+6(1)+16) =558 g